OpenStudy (anonymous):

What is the standard form of the equation that has a slope of 2/3 and passes through the point (1,0)

OpenStudy (anonymous):

One second. :)

OpenStudy (anonymous):

-2x + 3y = -2

OpenStudy (anonymous):

@DNADRE Is this what you are looking for?

OpenStudy (anonymous):

Use Point-Slope Form \[y - y_{1} = m(x - x_{1})\]\[y - 0 = \frac23(x - 1)\]\[y = \frac{2}{3}x - \frac{2}{3}\]\[-3y = -2x + 2\]\[2x - 3y = 2\]

OpenStudy (anonymous):

I just graphed the equation. And @Calcmathlete that is the same answer as mine, just multiplied by -1. I didn't think about using the point-slope form. Good job. :)